# Constructions of derived equivalent hyper-Kähler fourfolds

### v4

@@ 1. Introduction
Recall that, by considering equivalences of Hilbert schemes of points on $K3$ surfaces induced by equivalences between those $K3$ surfaces, non-isomorphic and derived equivalent $K3^{[n]}$-type manifolds with Picard group of rank $\geq 2$ were found in [P, MMY]. Moreover, examples of twisted derived equivalent $K3^{[2]}$-type manifolds that are moduli spaces of sheaves were described in [S, ADM]. In those cases Problem 1.1 has a positive answer. Moreover, we shall see that $K3^{[2]}-$ type manifolds with Hodge isometric transcendental lattices are always twisted derived equivalent. On the other hand, there exist $K3^{[2]}$-type manifolds of higher Picard number which have isometric transcendental lattice, but are not derived equivalent (cf. [MM, Theorem 5.13]).
@@ Acknowledgement
We thank N. Addington, E. Markman, D. Mattei, M. Mauri and T. Wawak for helpful discussions and K. Grzelakowski and C. Tschanz for comments.
@@ 2.2. Deformation of a derived equivalence
Arguing as in the proof of [Mar2, Theorem 1.4] (cf. [Mar2, Theorem 1.6]) using [Mar1, Lemma 8.3], we can find a special deformation $X_{t_{0}}$ of $X_{0}$ along a twistor path, such that $X_{t_{0}}$ is isomorphic to the square of a non-algebraic K3 surface and such that $E_{t_{0}}^{p}$ is poly-stable with respect to every Kähler class on $X_{t_{0}}$. It follows from [V1, Corollary 8.1] that the dimensions of cohomologies of poly-stable hyperholomorphic vector bundles are invariant along twistor lines. We infer that
@@ 3. Twisted derived equivalences of r-cyclic type
Theorem 3.1 (Theorem 1.5 [Mar2]).
If $f\colon H^{2}(X,{\mathbb{Q}})\to H^{2}(Y,{\mathbb{Q}})$ is a rational Hodge isometry of $r$-cyclic type for two hyper-Kähler manifolds $X$, $Y$ of $K3^{[n]}$-type then there exists a locally free possibly twisted sheaf on $X\times Y$ of rank $n!r^{n}$ that is a deformation along a twistor path of a BKR Fourier-Mukai kernel between two manifolds $S^{[n]}$ and $M^{[n]}$ for two derived equivalent K3 surfaces $S$ and $M$.
@@
For $X$ and $Y$ related by a rational Hodge isometry of r-cyclic type there exists a twisted derived equivalence of r-cyclic type: $D(X,\beta_{X})\simeq D(Y,\beta_{Y})$ for some Brauer classes $\beta_{X}\in H^{2}(\mathcal{O}^{\ast}_{X})$ and $\beta_{Y}\in H^{2}(\mathcal{O}^{\ast}_{Y})$.
@@ 3.2. Case of divisibility 1
identified via markings to $\frac{c_{1}(\mathcal{K}_{S^{[2]},M^{[2]}})}{2r^{2}}$. The latter, by [Mar2, Lemma 11.1] gives
(3.1) $\frac{c_{1}(\mathcal{K}_{S^{[2]},M^{[2]}})}{2r^{2}}=(\frac{e+rsf}{r}-\frac{\delta}{2},\frac{se+rf}{r}-\frac{\delta}{2})=(\frac{L_{r,s}}{r}+(\frac{l}{r}-\frac{1}{2})\delta,\frac{L^{\prime}_{r,s}}{r}+(\frac{l}{r}-\frac{1}{2})\delta).$
From Theorem 2.5 and (3.1) we infer that $\beta$ and $\beta^{\prime}$ are Brauer classes admitting lifts to $(\frac{l}{r}-\frac{1}{2})\delta_{X}$ and $(\frac{l}{r}-\frac{1}{2})\delta_{Y}$ respectively, where $\delta_{X}=\phi_{X}^{-1}(\delta)$ and $\delta_{Y}=\phi_{Y}^{-1}(\delta)$ In particular $\beta$ and $\beta^{\prime}$ are trivial when $\frac{2l}{r}$ is odd and equal to $[\frac{\delta_{X}}{2}]$ and $[\frac{\delta_{Y}}{2}]$ respectively when $\frac{2l}{r}$ is even. This concludes the proof. ∎
@@ 3.3. Case of divisbility 2
(3.2) $\frac{c_{1}(\mathcal{K}_{S^{[2]},M^{[2]}})}{2r^{2}}=(\frac{e+rsf}{r}-\frac{\delta}{2},\frac{se+rf}{r}-\frac{\delta}{2})=(\frac{\bar{L}_{r,s}}{2r}-\frac{l+r}{2r}\delta,\frac{\bar{L}^{\prime}_{r,s}}{2r}-\frac{l+r}{2r}\delta).$
@@ 6.1. Twisted fibrations
for each Brauer class $\beta$ and its $B$ lift $B$. Consequently we have exactly two torsors which are moduli spaces of twisted sheaves twisted with respect to the given Brauer class $\beta$.
@@ 6.3. Brauer class of universal families
and $R$ is another B lift of $T$. This prove the claim.
@@ 6.4. Duality between double EPW sextics and double EPW quartics
where $B^{\prime}$ is a B-lift of a corresponding Brauer class on $T$. Here $B\in H^{2}(S,\mathbb{Z})$ is a $B$-field lift of a Brauer class $\beta$ determined by the property that it has a B-lift with the properties $Bh=0$ $B^{2}=\frac{1}{2}$. It follows from Lemma 6.2 that $X_{0}$ (resp. $X_{1}$) can be seen as the relative $\operatorname{Pic}_{0}$ (resp. $\operatorname{Pic}_{1}$) over the twisted fibration ${\mathcal{C}}\to{\mathbb{P}}^{2}$ with Brauer class corresponding to $B$ through [Mar2, Lemma 7.3(4)].
@@ 6.5. Convolutions of two Fourier-Mukai kernels
The sheaf $F_{0}$ is a on rank $8$ vector bundle. Moreover, $F_{0}^{p}$ is slope-stable hyperholomorphic with respect to all Kähler classes on $X_{0}$ (resp. on $X_{0}^{\prime}$) for each $p\in X_{0}^{\prime}$ (resp. $p\in X_{0}$).
@@
The stability of $F_{0}$ follows from the fact that $X_{1}$ is irreducible. Indeed, it follows from [Mar1, p. 35] that it is enough to show that the restriction $F_{0}^{x}$ of $F_{0}$ is stable on $X_{0}\times\{x\}$ for a general $x\in X_{0}^{\prime}$ and on $\{y\}\times X_{0}^{\prime}$ for $y\in X_{0}^{\prime}$. Next, we find that the restriction of $F_{0}^{x}$ to a fiber of the Lagrangian fibration is a sum of different translations of the Theta divisor. It follows that a subbundle $E_{0}^{x}$ of $F_{x}^{0}$ that could destabilise $F_{0}^{x}$ restricts to partial sum of the line bundles before. Since each such line bundle corresponds to a point on $X_{1}$ through the Fourier-Mukai transform, the subbundle induces an analytic subvariety of $X_{1}$ that is open in the Zariski topology so gives a subvariety of $X_{1}$. So $E_{0}^{x}=F_{0}^{x}$ and $F_{0}^{x}$ is stable. ∎
@@ 7. Summary
The strategy described in the previous section can be applied in many cases. Let us describe several twisted derived equivalences of hyper-Kähler manifolds of Picard rank $\geq 2$. For a K3 surface $S$ of degree $2$ there are a lot of example of moduli spaces of twisted sheaves $M_{(0,h,0)}(S,B)$ that are hyper-Kähler fourfolds of $K3^{[2]}-$type with two Lagrangian fibrations that differ by the choice of a Brauer class $\beta\in H^{2}(\mathcal{O}_{S}^{\ast})$. More precisely, we choose a $k$-torsion Brauer class that admits a $B$ lift of class $B$ with

### v5

@@ 1. Introduction
Recall that, by considering equivalences of Hilbert schemes of points on $K3$ surfaces induced by equivalences between those $K3$ surfaces, non-isomorphic and derived equivalent $K3^{[n]}$-type manifolds with Picard group of rank $\geq 2$ were found in [P, MMY]. Moreover, examples of twisted derived equivalent $K3^{[2]}$-type manifolds that are moduli spaces of sheaves were described in [S, ADM]. In those cases Problem 1.1 has a positive answer. Moreover, we shall see that $K3^{[2]}-$ type manifolds with Hodge isometric transcendental lattices are always twisted derived equivalent. On the other hand, there exist $K3^{[n]}$-type manifolds (for each $n>2$) of Picard number two which have isometric transcendental lattice, but are not derived equivalent (cf. [MM]).
@@
$\begin{cases}2^{\tau(d)}\text{ when }\rm{div}(L_{X})=1,\\ 2^{\tau(d)-1}\text{ when }\rm{div}(L_{X})=2,\end{cases}$
where $\tau(d)$ is the number of different prime divisors of $d$ and $\tau(1)=1$.
@@ Acknowledgement
We thank N. Addington, E. Markman, D. Mattei, M. Mauri, T. Wawak and Ruxuan Zhang for helpful discussions and K. Grzelakowski and C. Tschanz for comments.
@@ 2.2. Deformation of a derived equivalence
Arguing as in the proof of [Mar2, Theorem 1.4] (cf. [Mar2, Theorem 1.6]) using [Mar1, Lemma 8.3], we can find a special deformation $X_{t_{0}}$ of $X_{0}$ along a twistor path, such that $X_{t_{0}}$ is isomorphic to the square of a non-algebraic K3 surface and such that $E_{t_{0}}^{p}$ is poly-stable with respect to every Kähler class on $X_{t_{0}}$. It follows from [V1, Corollary 8.1] that the dimensions of cohomologies of poly-stable hyperholomorphic vector bundles are invariant along twistor lines. We infer using the fact that $E_{t}^{p}$ is locally free so $\mathcal{E}xt^{1}(E_{t}^{p},E_{t}^{p})=0$ that
@@ 3. Twisted derived equivalences of r-cyclic type
Theorem 3.1 (Corollary 8.5 [Mar1]).
If $f\colon H^{2}(X,{\mathbb{Q}})\to H^{2}(Y,{\mathbb{Q}})$ is a rational Hodge isometry of $r$-cyclic type for two hyper-Kähler manifolds $X$, $Y$ of $K3^{[n]}$-type for which there exists a Kähler class $\kappa\in H^{2}(X,{\mathbb{Q}})$ such that $f(\kappa)$ is a Kähler class, then there exists a locally free possibly twisted sheaf on $X\times Y$ of rank $n!r^{n}$ that is a deformation along a twistor path of a BKR Fourier-Mukai kernel between two manifolds $S^{[n]}$ and $M^{[n]}$ for two derived equivalent K3 surfaces $S$ and $M$.
@@
For $X$ and $Y$ hyper-Kähler manifolds of $K3^{[n]}$-type of Picard number one related by a rational Hodge isometry of r-cyclic type there exists a twisted derived equivalence of r-cyclic type: $D(X,\beta_{X})\simeq D(Y,\beta_{Y})$ for some Brauer classes $\beta_{X}\in H^{2}(\mathcal{O}^{\ast}_{X})$ and $\beta_{Y}\in H^{2}(\mathcal{O}^{\ast}_{Y})$.
@@ 3.2. Case of divisibility 1
identified via markings to $\frac{c_{1}(\mathcal{K}_{S^{[2]},M^{[2]}})}{2r^{2}}$. The latter, by [Mar2, Lemma 11.1] and [Mar1, Eq. (7.1)-(7.2)] gives
(3.1) $\frac{c_{1}(\mathcal{K}_{S^{[2]},M^{[2]}})}{2r^{2}}=(\frac{e+rsf}{r}-\frac{\delta}{2},\frac{se+rf}{r}+\frac{\delta}{2})=(\frac{L_{r,s,l}}{r}+(\frac{l}{r}-\frac{1}{2})\delta,\frac{L^{\prime}_{r,s,l}}{r}+(\frac{l}{r}+\frac{1}{2})\delta).$
From Theorem 2.5 and (3.1) we infer that $\beta$ and $\beta^{\prime}$ are Brauer classes admitting lifts to $(\frac{l}{r}-\frac{1}{2})\delta_{X}$ and $(\frac{l}{r}+\frac{1}{2})\delta_{Y}$ respectively, where $\delta_{X}=\phi_{X}^{-1}(\delta)$ and $\delta_{Y}=\phi_{Y}^{-1}(\delta)$. In particular $\beta$ and $\beta^{\prime}$ are trivial when $\frac{2l}{r}$ is odd and equal to $[\frac{\delta_{X}}{2}]$ and $[\frac{\delta_{Y}}{2}]$ respectively when $\frac{2l}{r}$ is even. This concludes the proof. ∎
@@ 3.3. Case of divisbility 2
(3.2) $\frac{c_{1}(\mathcal{K}_{S^{[2]},M^{[2]}})}{2r^{2}}=(\frac{e+rsf}{r}-\frac{\delta}{2},\frac{se+rf}{r}+\frac{\delta}{2})=(\frac{\bar{L}_{r,s,l}}{2r}-\frac{l+r}{2r}\delta,\frac{\bar{L}^{\prime}_{r,s,l}}{2r}-\frac{l-r}{2r}\delta).$
@@ 6.1. Twisted fibrations
for each Brauer class $\beta$ and its $B$-lift $B$. Consequently we have exactly two torsors which are moduli spaces of twisted sheaves twisted with respect to the given Brauer class $\beta$.
@@ 6.3. Brauer class of universal families
and $R$ is another $B$-lift of $T$. This prove the claim.
@@ 6.4. Duality between double EPW sextics and double EPW quartics
where $B^{\prime}$ is a $B$-lift of a corresponding Brauer class on $T$. Here $B\in H^{2}(S,\mathbb{Z})$ is a $B$-field lift of a Brauer class $\beta$ determined by the property that it has a $B$-lift with the properties $Bh=0$, $B^{2}=0$. It follows from Lemma 6.2 that $X_{0}$ (resp. $X_{1}$) can be seen as the relative $\operatorname{Pic}_{0}$ (resp. $\operatorname{Pic}_{1}$) over the twisted fibration ${\mathcal{C}}\to{\mathbb{P}}^{2}$ with Brauer class corresponding to $B$ through [Mar2, Lemma 7.3(4)].
@@ 6.5. Convolutions of two Fourier-Mukai kernels
The sheaf $F_{0}$ is a on rank $8$ vector bundle. Moreover, $F_{0}^{p}$ is slope-stable and $\mathcal{E}nd(F_{0}^{p})$ hyperholomorphic with respect to all Kähler classes on $X_{0}$ (resp. on $X_{0}^{\prime}$) for each $p\in X_{0}^{\prime}$ (resp. $p\in X_{0}$).
@@
The stability of $F_{0}$ follows from the fact that $X_{1}$ is irreducible. Indeed, it follows from [Mar1, p. 35] that it is enough to show that the restriction $F_{0}^{x}$ of $F_{0}$ is stable on $X_{0}\times\{x\}$ for a general $x\in X_{0}^{\prime}$ and on $\{y\}\times X_{0}^{\prime}$ for $y\in X_{0}^{\prime}$. Next, we find that the restriction of $F_{0}^{x}$ to a fiber of the Lagrangian fibration is a sum of different translations of the Theta divisor. It follows that a subbundle $E_{0}^{x}$ of $F_{x}^{0}$ that could destabilise $F_{0}^{x}$ restricts to partial sum of the line bundles before. Since each such line bundle corresponds to a point on $X_{1}$ through the Fourier-Mukai transform, the subbundle induces an analytic subvariety of $X_{1}$ that is open in the Zariski topology so gives a subvariety of $X_{1}$. So $E_{0}^{x}=F_{0}^{x}$ and $F_{0}^{x}$ is stable.
We conclude from [Mar1, Section 5.6] that $\mathcal{E}nd(F_{0}^{x})$ is hyperholomorphic. ∎
@@ 7. Summary
The strategy described in the previous section can be applied in many cases. Let us describe several twisted derived equivalences of hyper-Kähler manifolds of Picard rank $\geq 2$. For a K3 surface $S$ of degree $2$ there are lots of examples of moduli spaces of twisted sheaves $M_{(0,h,0)}(S,B)$ that are hyper-Kähler fourfolds of $K3^{[2]}-$type with two Lagrangian fibrations that differ by the choice of a Brauer class $\beta\in H^{2}(\mathcal{O}_{S}^{\ast})$. More precisely, we choose a $k$-torsion Brauer class that admits a $B$-lift of class $B$ with